# Know “No” Better: A Data-Driven Approach for Enhancing Negation Awareness in CLIP

### v2

@@ 1 Introduction
Recent advances in vision-language models (VLMs) [19, 16, 51, 20] have demonstrated significant capabilities in integrating visual and linguistic information, achieving notable performance in multimodal tasks such as text-to-image (T2I) generation and referring image segmentation [39, 2]. Among these models, CLIP [35] has emerged as particularly influential, serving as the foundation for numerous subsequent models [38, 25, 28, 36]. The effectiveness of these models, however, is inherently constrained by the capabilities of the CLIP encoder, underscoring the importance of its robustness.
@@
Our experiments demonstrate that NegationCLIP, fine-tuned using data generated by our proposed framework, achieves superior negation understanding on both existing and newly developed NegRefCOCOg benchmarks while maintaining strong general task performance. Additionally, NegationCLIP proves adaptable across various multimodal tasks. Notably, replacing the text encoder in T2I generation models with that of NegationCLIP enhances negation comprehension—a capability often lacking in original T2I models (using the original CLIP encoder). Furthermore, our NegationCLIP’s text encoder enables improved performance in referring image segmentation for prompts containing negation, underscoring its scalability as a more contextually aware and flexible text encoder. Know “No” Better: A Data-Driven Approach for Enhancing Negation Awareness in CLIP illustrates the versatility of NegationCLIP across diverse multimodal tasks involving negation. The key contributions of our study are summarized as follows:
@@ Appendix D NegRefCOCOg Benchmark
We obtain 440 triplets of ($T$, $P^{+}$, $P^{-}$) through this process, where $P^{+}$ is well-aligned with $T$ and $P^{-}$ serves as challenging hard negative for evaluation. These 440 samples form the NegRefCOCOg benchmark, which we use to evaluate the ability of models to handle negation comprehensively and accurately.
@@
Appendix E Experiment Details
E.1 CelebA Classification
@@
E.2 Fine-Tuning Configuration
@@ E.3 T2I Generation
E.4 Referring Image Segmentation
@@
Appendix F Expanded Discussion on T2I Generation
F.1 Methods
F.2 Evaluation
Table 12: Results on LMD Negation benchmark.
Method LMD Negation
SD 20%
SD w/ NegationCLIP text encoder 94%
SD + LMD [22] 100%
Appendix G Additional Qualitative Results

### v3

@@ 1 Introduction
Recent advances in vision-language models (VLMs) [22, 16, 53, 23, 55] have demonstrated significant capabilities in integrating visual and linguistic information, achieving notable performance in multimodal tasks such as text-to-image (T2I) generation [18, 41, 3] and referring image segmentation [31, 19, 20]. Among these models, CLIP [38] has emerged as particularly influential, serving as the foundation for numerous subsequent models [41, 28, 31, 39]. The effectiveness of these models, however, is inherently constrained by the capabilities of the CLIP encoder, underscoring the importance of its robustness.
@@
Our experiments demonstrate that NegationCLIP, fine-tuned using data generated by our proposed framework, achieves superior negation understanding on both existing and newly developed NegRefCOCOg benchmarks while maintaining strong general task performance. Additionally, NegationCLIP proves adaptable across various multimodal tasks. Notably, replacing the text encoder in T2I generation models with that of NegationCLIP enhances negation comprehension—a capability often lacking in original T2I models (using the original CLIP encoder). Furthermore, our NegationCLIP’s text encoder enables improved performance in referring image segmentation for prompts containing negation, underscoring its scalability as a more contextually aware and flexible text encoder. Fig. 1 illustrates the versatility of NegationCLIP across diverse multimodal tasks involving negation. The key contributions of our study are summarized as follows:
@@ 7 Conclusion
Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2022R1A3B1077720, 2022R1A5A7083908, RS-2025-00555943), Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [NO. RS-2021-II211343, RS-2022-II220959, Artificial Intelligence Graduate School Program (Seoul National University), No.RS-2025-02263754, Human-Centric Embodied AI Agents with Autonomous Decision-Making], the BK21 FOUR program of the Education and Research Program for Future ICT Pioneers, Seoul National University in 2025, and Samsung Electronics (IO221213-04119-01).
@@ Appendix D NegRefCOCOg Benchmark
We obtain 440 triplets of ($T$, $^{+}, PP^{-}$) through this process, where $P^{+}$ is well-aligned with $T$ and $P^{-}$ serves as challenging hard negative for evaluation. These 440 samples form the NegRefCOCOg benchmark, which we use to evaluate the ability of models to handle negation comprehensively and accurately.
@@
Appendix E Results on Additional Models
Our method is model-agnostic and applies to SigLIP [55], showing consistent gains on negation benchmarks while preserving general performance, as shown in Tab. 10.
Table 10: Comparison of model performance on negation and general benchmarks across different architectures of SigLIP [55].
Model Arch. Negation Benchmarks General Benchmarks
VALSE NegRefCOCOg ImageNet COCO
SigLIP base 67.98 61.59 75.69 72.10
Neg.SigLIP 77.15 68.18 75.10 77.30
SigLIP large 73.03 62.05 79.73 75.20
Neg.SigLIP 79.40 67.73 79.76 80.10
SigLIP so400m 75.47 63.18 82.26 76.30
Neg.SigLIP 84.27 67.05 81.68 81.70
Appendix F Expanded Discussion on T2I Generation
F.1 Methods
F.2 Evaluation
Table 11: Results on LMD Negation benchmark.
Method LMD Negation
SD 20%
SD w/ NegationCLIP text encoder 94%
SD + LMD [25] 100%
Appendix G Experiment Details
G.1 CelebA Classification
@@
G.2 Fine-Tuning Configuration
@@ G.3 T2I Generation
G.4 Referring Image Segmentation
@@
Appendix H Additional Qualitative Results